I have these two equations: 1/3x+y=1 and -4x+y=1
I have to solve it using elimination. What do I do to isolate a single variable?

Answer by edjones(8007)

1/3x+y=1
-4x+y=1
.
1/3x=-4x Since either number added to y equals 1.
1=-12x^2
x^2=-1/12
x=+-sqrt(1/12)i
=+-sqrt(1/3)/2 *i
x=sqrt(1/3)/2 *i, -sqrt(1/3)/2 *i
.
Ed